# Symmetry-Informed Governing Equation Discovery

### v1

@@ C.2 Symmetry Regularization
Table 4: Equation discovery statistics on Lotka-Volterra system (15) at noise level $\sigma_{R}=99\%$ and Sel’Kov oscillator (16) at noise level $\sigma_{R}=20\%$. The RMSE is scaled by ×$\times 10^{-1}$ for (15) and ×$\times 10^{-2}$ for (16). The success probability is computed from 50 runs for sparse regression-based algorithms and 10 runs for genetic programming-based algorithms. The success probabilities of recovering individual equations (Eq. 1 & Eq. 2) and simultaneously recovering both equations (All) are reported. The RMSE (all) refers to the parameter estimation error over all runs. The RMSE (successful) refers to the parameter estimation error over successful runs, which is missing for algorithms with zero success probability.
System Method Success prob. RMSE (successful) RMSE (all)
Eq. 1 Eq. 2 All Eq. 1 Eq. 2 All Eq. 1 Eq. 2
GP 1.0 0.0 0.0 2.44 (0.89) N/A N/A 2.44 (0.89) 13.20 (3.20)
D-CODE 0.0 0.0 0.0 N/A N/A N/A 10.38 (0.11) 9.08 (1.88)
Sel’Kov (16) SINDy 0.30 0.56 0.14 0.87 (0.15) 0.32 (0.13) 0.67 (0.10) 15.86 (17.52) 12.71 (18.32)
GP 0.0 0.0 0.0 N/A N/A N/A 7.57 (5.46) 14.88 (23.50)
D-CODE 0.0 0.4 0.0 N/A 0.27 (0.11) N/A 14.37 (15.66) 8.90 (16.84)
In addition to Table 2 in the main text, Table 4 shows the parameter estimation RMSE over (1) successful runs and (2) all runs. Similar to the previous observations, our methods based on symmetry regularization have the lowest parameter estimation errors averaging over all random experiments. Also, Weak SINDy (WSINDy) has the most accurate parameter estimation conditioning on the successful discovery of the correct equation form. The potential reason for this phenomenon has been discussed in Section C.1.
@@
Figure 5 and Figure 6 show the error growth when using the discovered ODEs to simulate future dynamics. Some algorithms are not included in the plot because the simulation from the learned equation quickly diverges and the error grows to infinity. For the Lotka-Volterra system in Figure 5, we observe that the prediction error grows slowest when using the discovered equations from our methods based on symmetry regularizations. Among all the methods, genetic programming with symmetry regularization has the most accurate prediction.
@@
where $g=\exp(\epsilon v)$ is computed for some fixed scale parameter $\epsilon$.
@@
Besides, (36) and (37) have a specific advantage that any computation related to group actions is independent of the function $\bm{h}$ which is constantly updated throughout training. Thus, we can compute the (infinitesimal) group actions and their Jacobians in advance over the entire dataset. This provides better compatibility between the use of symmetry and existing tools for equation discovery. For example, in PySR framework (Cranmer, 2023), it is difficult to compute group transformations $g\cdot\bm{x}$ based on an external neural network module during the equation seearch. We overcome this challenge by pre-computing the transformations and using the transformed data as auxiliary inputs.
@@ D.2.2 Genetic Programming
We incorporate our reversed symmetry regularization to an existing genetic programming package for discovering equations, PySR (Cranmer, 2023). The “GP” baseline in the tables also refers to the implementation of genetic programming with PySR. On the other hand, we run the experiments on D-CODE with their official codebase  [^4] https://github.com/ZhaozhiQIAN/D-CODE-ICLR-2022. This codebase is developed based on another genetic programming package, gplearn. We should note that the performance difference between our EquivGP and D-CODE can be partially caused by their different backends.

### v2

@@ C.1 Equivariant SINDy for Linear Symmetries
C.1.2 A Higher-Dimensional System
Our method can also be applied to higher-dimensional dynamical systems with symmetries. To demonstrate this, we consider the SEIR equations Abou-Ismail (2020) which model the evolution of pandemics by the number of susceptible (S), exposed (E), infected (I), and recovered (R) individuals. It is a 4-dimensional ODE system with quadratic terms. We use the following equations as the target of equation discovery:
[…]
Compared to the original formulation, we add two constant terms to the first and the fourth equation to account for the re-infection of patients due to waning immunity, mutation of the pathogen, etc. During data generation, we add a moderate $5\%$ noise to all trajectories.
We assume an intuitive symmetry of scaling the number of recovered individuals proportional to the total population: $v=(S+E+I+R)\partial_{R}$. We solve the constraint with regard to this symmetry as in Section 4.1. Figure 5 shows the reduced parameter space (from 60D to 34D).
Figure 5: In the SEIR model, we search the equations of 4 variables with up to quadratic terms, leading to a 4$\times$15=60D parameter space. The symmetry $v=(S+E+I+R)\partial_{R}$ reduces it to 34D.
Table 4 shows the discovery results of SINDy and our EquivSINDy-c in terms of success probabilities and parameter estimation error when successful. It turns out that SINDy struggles to find the correct equations and never manages to correctly identify all the equations simultaneously. In comparison, by reducing the parameter space through symmetry constraint, our method greatly increases the chance of successful discovery.
Table 4: Equation discovery on SEIR epidemic model (4D). Results are evaluated on 50 random runs.
Dynamics Method Success prob. RMSE (successful)
Individual eqs Joint Individual eqs Joint
SEIR SINDy 0.02/0.02/0.20/0.10 0.00 1.57/0.39/1.42/1.54 -
EquivSINDy-c 0.26/0.98/0.84/0.64 0.14 1.22/0.55/1.86/1.62 1.57 (0.47)
@@ C.2 Symmetry Regularization
In addition to Table 2 in the main text, Figure 6 and Figure 7 show the error growth when using the discovered ODEs to simulate future dynamics. Some algorithms are not included in the plot because the simulation from the learned equation quickly diverges and the error grows to infinity. For the Lotka-Volterra system in Figure 6, we observe that the prediction error grows slowest when using the discovered equations from our methods based on symmetry regularizations. Among all the methods, genetic programming with symmetry regularization has the most accurate prediction.
@@
where $g=\exp(\epsilon v)$ is computed for some fixed scale parameter $\epsilon$. In other words, the summation is always performed over the learned infinitesimal generators ($v\in B(\mathfrak{g})$ and the group elements $g$ in the above formulas are dependent on $v$.
We also note that the denominator of (36) has a somewhat different form from others, i.e. the difference between $\bm{f}(g\bm{x})$ and $\bm{f}(\bm{x})$. If we use $\bm{f}(g\bm{x})$ or $\bm{f}(\bm{x})$ instead, the scale of the denominator would also depend on $\bm{x}$. As a result, regions closer to the origin in the state space have larger weights (smaller denominators) during loss computation. It’s helpful to consider what would happen in the limit of $\epsilon\to 0,\tau\to 0$. The formulas inside the norms in the numerator and denominator in (35) both become $\bm{O}(\epsilon\tau)$, while $\bm{f}(g\bm{x})=\bm{f}(\exp(\epsilon v)\bm{x})=\bm{f}(\bm{x}+\bm{O}(\epsilon))=\bm{x}+\bm{O}(\epsilon)+\bm{O}(\tau)+\bm{O}(\epsilon\tau)$, which does not match the numerator.
@@
Besides, (37) and (38) have a specific advantage that any computation related to group actions is independent of the function $\bm{h}$ which is constantly updated throughout training. Thus, we can compute the (infinitesimal) group actions and their Jacobians in advance over the entire dataset. This provides better compatibility between the use of symmetry and existing tools for equation discovery. For example, in PySR framework (Cranmer, 2023), it is difficult to compute group transformations $g\cdot\bm{x}$ based on an external neural network module during the equation search. We overcome this challenge by pre-computing the transformations and using the transformed data as auxiliary inputs.
@@ C.3 Equivariant SINDy in Latent Space
C.4 Comparison with Neural-ODE-Based Methods
We compare our method with the Time-Reversal Symmetry ODE Network (TRS-ODEN) (Huh et al., 2020), which uses the discrete symmetry of time-reversal to improve dynamics learning. TRS-ODEN aims to learn the physical dynamics with black-box neural networks and adds a loss term to encourage time-reversal symmetry. We evaluate TRS-ODEN as well as other baselines in Huh et al. (2020), i.e. Neural ODE (ODEN) and Hamiltonian Neural ODE (HODEN), on the damped oscillator system (5.1) and the Lotka-Volterra system (15) in terms of the long-term prediction error. Other metrics, including success probability and parameter estimation error, are not compared because Neural ODE does not learn a symbolic equation.
Figure 9: Long-term prediction errors of our method & Neural-ODE-based methods on the irreversible damped oscillator (left) and the reversible Lotka-Volterra system (right).
Figure 9 plots the long-term prediction error against the prediction horizon. In the damped oscillator, because it is irreversible and not conservative, TRS-ODEN and HODEN do not help. In the reversible Hamiltonian system of Lotka-Volterra equations, TRS-ODEN and HODEN improve upon the baseline Neural ODE, but have slightly higher errors than our discovered equations with Lie symmetry regularization. Notably, in the Lotka-Volterra system, our method does not rely on any prior knowledge (e.g. time-reversal symmetry or conservation of energy) but still achieves lower prediction error. To conclude, these results demonstrate that our method can be universally applied to various dynamical systems and achieve superior performance.
@@ D.2.2 Genetic Programming
We incorporate our symmetry regularization (37) to an existing genetic programming package for discovering equations, PySR (Cranmer, 2023). The “GP” baseline in the tables also refers to the implementation of genetic programming with PySR. On the other hand, we run the experiments on D-CODE with their official codebase  [^4] https://github.com/ZhaozhiQIAN/D-CODE-ICLR-2022. This codebase is developed based on another genetic programming package, gplearn. We should note that the performance difference between our EquivGP and D-CODE can be partially caused by their different backends.